Problem 5-14 Calculating Interest Rates [LO3] Assume that in 2010, a gold dollar minted in 1886 sold for $134,000. For this to have been true, what rate of return did this coin return for the lucky numismatist? (Do not round intermediate calculations and round your final answer to 2 decimal places. (e.g., 32.16)) Rate of return
Answer #1

Value in 1886 = $1

Value in 2010 = $134,000

Number of years = 2010-1886 = 124 years

Future value = Present value * (1 + r)n

$134,000 = $1*(1+r)124

1+r = 1340001/24

r = 1.6354 - 1 = 0.6354

Rate of return = 63.54%